# Quantifying Uncertainty: All We Need is the Bootstrap?

### v1

@@ Abstract
Standard errors, confidence intervals, hypothesis tests, and other quantifications of uncertainty are essential to statistical practice. However, they feature a plethora of different methods, mathematical formulas, and concepts. Could we not just replace them all with the general and relatively easy-to-understand non-parametric bootstrap? We contribute to answering this question with a review of related work and a simulation study of one- and two-sided confidence intervals over several different sample sizes, confidence levels, data generating processes, and functionals. Results show that double bootstrap is the overall best method and a viable alternative to typically used approaches in all but the smallest sample sizes.
statistics,
inference,
standard errors,
confidence intervals,
simulation study,
@@
and
@@ 1 Introduction
University curricula in the social sciences, medicine, life sciences, and other fields that heavily utilize statistical methodology typically include only one or two applied statistics courses. However, it is these practitioners, rather than professional statisticians, who end up conducting most of the statistical analyses. The disparity between the level of training received and the demand for statistical analysis leads to an over-reliance on rote learning and formulaic application of statistical methods, which likely contributes to the replication crisis in science. As such, investigating all opportunities for simplifying aspects of current statistical practice is worthwhile. Such simplification could not only improve understanding and reduce mistakes but also make room in curricula for other parts of statistical methodology.
In this paper we focus on quantification of uncertainty. Standard errors, confidence intervals, and hypothesis tests are essential to statistical practice, but this part of statistical methodology typically features advanced concepts, such as test statistics and sampling distributions, and a plethora of different methods. One methods stands out, however. The bootstrap has several advantages over other methods for quantifying uncertainty. It is conceptually simple, it illustrates the central role of sampling in statistics, we can work directly with estimates and their distributions, and it can be applied to most practical tasks without having to learn new concepts or mathematical formulas. This gives bootstrapping high pedagogical value [25] and makes it an ideal candidate for a one-size-fits-all approach to quantifying uncertainty for practitioners that will receive a limited amount of statistical training.
Historically, bootstrapping was underutilized because it was computationally prohibitive. Today, that is no longer the case, yet introductory applied statistics books and courses where the bootstrap takes central stage are still rare. From our experience, this is mostly due to the convenience of following established practices, which do have better instructional materials and software (something that bootstrapping lacked [25] and still does). However, evidence that bootstrapping is indeed a viable one-size-fits-all approach to quantifying uncertainty is also too sparse to be convincing.
@@ 1.1 Related work
Theoretical cases where bootstrap fails are pathological with limited implications for practice (see [7] and [10, Ch.2.6]). Large sample properties of common bootstrap approaches have also been established (see [7] for a summary). Unfortunately, large sample theory is not always a reliable predictor of finite sample performance, so we have to turn to empirical work.
@@
[16] normal 14 200 mean SE PB, SB (B = 128, 512) SB better than PB
[42] normal 20-100 1600 variance 90% CI BC, PB (B = 1000) bootstrap coverage below nominal
[19] normal, exponential 15 200 trimmed mean SE PB (B = 200) PB better than jackknife
[40] normal, non-normal 5-60 1000 mean 95%, 99% CI baseline; PB (B = 500) baseline better than PB
[44] normal 14-100 200 mean SE PB, SB (B = 200) SB better than PB
[14] exponential 5 1 corr 95% CI BC, BCa, PB (B = ?) BCa good; PB and BC poor
bivariate non-normal 5 1 ratio 95% CI BC, BCa, PB (B = ?) BCa good; PB and BC poor
normal 8 1 mean 95% CI B-t, BC, PB (B = ?) BC and PB perform similarly, B-t is worse
[17] normal 100 100 mean several CI baseline; PB, SB (B = 2000) PB similar to baseline; SB better than PB
[36] chi-squared 20 1000 corr 90% CI baseline; hybrid, B-t, BC, BCa, PB (B = 1000) hybrid, B-t, and PB perform similarly; BC and BCa have lower than nominal coverage
[15] (see paper) 13, 15 500 regression 90% CI baseline; BC, BCa, PB (B = 500) baseline better than bootstrap; bootstrap methods similar to each other
[47] normal, exponential 10, 20 1000 corr several CI PB (B = 50000) PB is biased
[37] normal, exponential, log-normal, mixture of normal, student, triangular, uniform 3-20 1000 corr 95% CI baseline; B-t, BCa, PB (B = 1000) B-t performs well
[34] chi-squared, mixture of normal 20 10000 corr several CI baseline; B-t, PB (B = 1000) bootstrap similar to baseline; B-t is best
[43] normal, Poisson, t, Weibull 20 500 corr 90% CI baseline; B-t, BCa, DB, PB (B = 500) DB and B-t perform best
[46] normal, exponential 5-25 1000 corr 95% CI baseline; DB, PB (B = 1000) DB as good or better than baseline, except for $n=5$; PB is worst
[30] normal, exponential, beta, gamma, t 10, 100 500 median SE PB, SB (B = 200) smoothing improves performance
[12] normal 20 1 mean 90% CI baseline; ABC, BCa, B-t (B = 2000) bootstrap better than baseline
[31] normal, folded normal, exponential, log-normal 15, 30 1600 corr 90% CI ABC, DB, DB-ABC, PB (B = 1000) DB better than ABC and calibrated ABC
[32] (see paper) 13, 15 500 regression 90% CI DB, PB (B = 1999) DB better than PB
[7] inverse exponential 20 10000 corr 99% CI baseline; B-t, BC, BCa, PB (B = 4999) baseline, B-t, and BCa perform similarly; BCa poor for small n
[18] normal 15 1 mean 90% CI baseline; ABC bootstrap better than baseline
[48] log-normal, gamma 10-50 10000 diff. in means 95% CI baseline; B-t, BCa (B = 1000) B-t performs well and better than BCa
[2] bivariate normal (censored) 25-400 2000 (see paper) 90%, 95% CI B-t, BCa, PB (B = 520) jackknife performs best; PB performs worst
[8] normal, log-normal, gamma, t, uniform 10-3600 1000-64000 variance several CI baseline; ABC, BC, BCa, PB (B = 1000, 16000) 2nd order accurate methods can converge slowly and perform worse than PB
[9] (see paper) 50-500 5000 regression 95% CI baseline; BC, PB (B = 2000) BC and PB perform well
[24] exponential, Pareto 10-25 2000 80%, 90% CI baseline; PB, DB (B = 699) DB is best
[28] (see paper) 50-500 5000 regression 95% CI baseline; BCa, PB (B = 10000) BCa and PB perform well
[29] (see paper) 10-50 1000 capture-recapture 95% CI baseline; B-t, PB (B = 100, 250) B-t performs best
[1] log-logistic with censored data 25-50 1000 (see paper) 90%, 95% CI baseline; B-t, DB, DB-t, PB (B = ?) PB poor; other bootstrap methods good and similar to each other
[38] normal, uniform, triangular, beta, Laplace, Pareto 50?-300? 1000 mean 95% CI BCa, PB (B = 2000) BCa slightly better than PB; good coverage except on Pareto
[25] normal, exponential 5-4000 10000 corr 95% CI baseline; B-t, PB, reverse PB (B = 10000) bootstrapped CI narrow, especially for small n; PB worse than baseline for small n; reverse PB poor; B-t best
[33] log-logistic with censored data 25-60 1000 (see paper) 90%, 95% CI baseline, JK; DB, PB (B = 1000) DB best; PB poor
[39] normal 10-100 10000 mean 95% CI baseline; B-n, B-t, BC, BCa, PB (B = 10000) baseline better than bootstrap on small n; BCa performs well
[21] (see paper) ? 1000 quantiles 95% CI modified B-n, BC, BCa, PB (B = 1999) bootstrap methods perform similarly
[26] log-normal 20-120 100 quantiles 95% CI baseline; BCa, PB (B = 10000) baseline better than bootstrap for small n
[11] 26 pairs 20-600 1000 corr 95% CI baseline; BCa, PB (B = 2000) BCa better than PB on small n; PB better than BCa for large n; poor coverage for uniform-chi-squared
Table 1 lists empirical work on bootstrap methods. Most of the early work was on Pearson correlation and the mean, while later work includes other functionals, mostly quantiles and regression parameters. We focus on non-parametric bootstrap methods and on methods that are most commonly taught and used for that task (from now on, baseline methods). For example, Fisher confidence intervals (CIs) for Pearson correlation and t-intervals for the mean.
Most but far from all related work includes such a baseline. Excluding an early study where the best bootstrap was the 1st-order accurate percentile bootstrap (PB) [40], there are only 3 studies where the baseline outperforms the best bootstrap method. Dorfman et al. [15] showed that Fieller and Taylor series-based intervals for elasticities and flexibilities outperform bias-corrected and accelerated (BCa), bias-corrected (BC), and PB, albeit modestly and on relatively small sample sizes ($n=13,15$). Puth et al. [39] show for Pearson correlation and the bivariate normal with three different correlation coefficients that Fisher intervals achieve coverage close to nominal 95%, while BCa achieves coverage around 93% and 94% for sample sizes 20 and 10, respectively. And Ialongo [27] showed for the log-normal distribution that the baseline attained nominal coverage at lower samples sizes than BCa, but only if log-normality was assumed.
To summarize, most related work is limited to a single functional, a single data generating process (DGP), and/or a single confidence level. The most commonly used method(s) are not always included as a baseline for comparison. And DB deserves more attention. So, while the results are encouraging, it is difficult to gauge the practical consequences of using only bootstrapping.
@@ 2 Simulation Study
We used $n_{\text{rep}}=10000$ replications for every experiment, to limit coverage standard error to $0.005$ in the worst case, and $B=\{10,100,1000\}$ bootstrap replications. More bootstrap replications is better, so we do not consider $B$ a dimension. We only report results for $B=1000$, but we would reach the same conclusions with $B=100$. For $B=10$ the performance of bootstrap methods is noticeably worse. Recommendations [8, 10, 20] and choices of $B$ in related work also suggest that $B=1000$ is enough.
Table 2: Mean KL coverage performance of bootstrap methods for one-sided CIs. The overall column is across all combinations, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined.
overall 4 8 16 32 64 128 256 corr mean median Q(0.05) Q(0.95) std
B-n $\underline{0.082}$ $0.322$ $\underline{0.180}$ $\underline{0.091}$ $\underline{0.038}$ $0.021$ $0.014$ $0.010$ $0.006$ $0.030$ $0.004$ $\underline{0.024}$ $\underline{0.145}$ $0.208$
B-t $0.088$ $\underline{0.199}$ $0.280$ $0.112$ $0.051$ $0.024$ $0.020$ $0.014$ $0.008$ $0.015$ $0.012$ $0.078$ $0.264$ $\underline{0.084}$
BB $0.117$ $0.253$ $0.194$ $0.227$ $0.106$ $0.047$ $0.034$ $0.023$ $0.017$ $0.034$ $0.045$ $0.159$ $0.234$ $0.145$
SB $0.124$ $0.459$ $0.278$ $0.160$ $0.063$ $0.028$ $0.018$ $0.011$ $0.002$ $0.026$ $0.003$ $0.064$ $0.230$ $0.308$
DB $0.140$ $0.544$ $0.348$ $0.200$ $0.060$ $\underline{0.008}$ $\underline{0.005}$ $\underline{0.003}$ $\underline{0.000}$ $\underline{0.010}$ $\underline{0.002}$ $0.125$ $0.401$ $0.188$
BCa $0.159$ $0.600$ $0.335$ $0.242$ $0.096$ $0.018$ $0.010$ $0.005$ $0.001$ $0.023$ $0.017$ $0.201$ $0.379$ $0.218$
BC $0.162$ $0.607$ $0.350$ $0.234$ $0.096$ $0.022$ $0.013$ $0.008$ $0.001$ $0.026$ $0.017$ $0.169$ $0.387$ $0.246$
PB $0.165$ $0.620$ $0.378$ $0.229$ $0.076$ $0.029$ $0.019$ $0.012$ $0.002$ $0.028$ $0.003$ $0.109$ $0.377$ $0.330$
@@ 2.1 Methods
With a one-size-fits-all method in mind, we include only non-parametric bootstrap methods that do not require user input: basic bootstrap (BB[20, chap. 13.4], standard bootstrap (B-n[20, chap. 13], B-t [20, chap. 14], BC [20, chap. 14], BCa [20, chap. 14], DB [22, chap. 3.11], PB [20, chap. 13], and smoothed bootstrap (SB[10].
@@
As a baseline for comparison, we include the CIs produced by the following commonly used methods: for the mean we use the t-based CIs from the commonly used t-test (t-test). For the median we use CIs from the Wilcoxon signed rank test (wilcoxon[3]. For standard deviation, we use the chi-squared CIs (chi-sq[45]. For Pearson correlation we use Fisher CIs (fisher) [4]. For quantiles we use parametric CIs based on normal assumption (q-par), non-parametric CI (q-nonpar) (see [27] for both), and the Maritz-Jarrett method (m-j) [35].

### v2

@@ Abstract
Quantifying uncertainty through standard errors, confidence intervals, hypothesis tests, and related measures is a fundamental aspect of statistical practice. However, these techniques involve a variety of methods, mathematical formulas, and underlying concepts, which can be complex. Could the non-parametric bootstrap, known for its simplicity and general applicability, serve as a universal alternative? In this study, we address this question through a review of existing literature and a simulation analysis of one- and two-sided confidence intervals across varying sample sizes, confidence levels, data-generating processes, and statistical functionals. Our findings indicate that the double bootstrap consistently performs best and is a promising alternative to traditional methods used for common statistical tasks. These results suggest that the bootstrap, particularly the double bootstrap, could simplify statistical education and practice without compromising effectiveness.
@@
Statistics; inference; standard errors; confidence intervals; simulation study
@@ 1 Introduction
University curricula in fields such as the social sciences, medicine, and life sciences, which heavily rely on statistical methodology, typically include only one or two applied statistics courses. However, it is practitioners in these fields, rather than professional statisticians, who perform the majority of statistical analyses. The mismatch between the level of training provided and the practical demand for statistical analysis often results in an over-reliance on rote memorization and formulaic application of methods, contributing to challenges such as the replication crisis in science. This underscores the importance of exploring ways to simplify current statistical practices. Simplification could not only enhance comprehension and reduce errors but also create opportunities to incorporate other aspects of statistical methodology into curricula.
In this paper, we focus on the quantification of uncertainty. Standard errors, confidence intervals, and hypothesis tests are integral components of statistical practice, yet they typically involve advanced concepts—such as test statistics and sampling distributions—and encompass a wide array of methods. Among these, one method stands out: the bootstrap. The bootstrap offers several advantages over traditional techniques for quantifying uncertainty. It is conceptually straightforward, reinforces the fundamental role of sampling in statistics, allows direct interaction with estimates and their distributions, and can be applied to a wide range of practical tasks without requiring the mastery of new concepts or complex mathematical formulas. These characteristics give the bootstrap significant pedagogical value [1], positioning it as a strong candidate for a one-size-fits-all approach to quantifying uncertainty, particularly for practitioners with limited statistical training.
Historically, the use of the bootstrap was constrained by computational limitations. With advancements in computing power, this is no longer a concern, yet introductory applied statistics textbooks and courses rarely place the bootstrap at the forefront. In our experience, this is largely due to the inertia of established practices, which are supported by more extensive instructional resources and software tools—resources that bootstrapping has historically lacked [1], and to some extent still does. Moreover, there is currently insufficient empirical evidence to convincingly establish the bootstrap as a comprehensive, one-size-fits-all solution for quantifying uncertainty.
@@ 1.1 Related work
Theoretical cases where bootstrap fails are pathological with limited implications for practice (see [2] and [3, Ch.2.6]). Large sample properties of common bootstrap approaches have also been established (see [2] for a summary). Unfortunately, large sample theory is not always a reliable predictor of finite sample performance, so we have to turn to empirical work.
Tables 1.1 and 1.1 summarize empirical studies on bootstrap methods. Early research primarily focused on the Pearson correlation and the sample mean, with subsequent work extending to other functionals, particularly quantiles and regression parameters. In this review, we concentrate on non-parametric bootstrap techniques and methods that are widely taught and applied in practice, referred to hereafter as baseline methods. These include, for example, Fisher confidence intervals (CIs) for the Pearson correlation and t-intervals for the sample mean.
While most, though not all, related studies include a baseline comparison, there are only three instances—aside from an early study where the 1st-order accurate percentile bootstrap (PB) was the best-performing method [4]—in which the baseline outperforms the top-performing bootstrap approach. For example, [5] demonstrated that Fieller and Taylor series-based confidence intervals for elasticities and flexibilities marginally outperformed the bias-corrected and accelerated (BCa), bias-corrected (BC), and PB methods, though this result was observed with relatively small sample sizes ($n=13,15$). Similarly, [6] showed that for Pearson correlation under a bivariate normal distribution with varying correlation coefficients, Fisher intervals achieved coverage close to the nominal 95%, while BCa intervals yielded coverage rates around 93% and 94% for sample sizes of 20 and 10, respectively. Furthermore, [7] found that for the log-normal distribution, baseline methods attained nominal coverage with smaller sample sizes than BCa, though this was contingent on assuming log-normality.
In summary, much of the related research is confined to a single functional, a single data-generating process (DGP), and/or a single confidence level. Additionally, the most commonly used methods are not always included as baselines for comparison. DB also warrants greater attention. While the findings are promising, it remains challenging to fully assess the practical implications of relying exclusively on bootstrap methods.
@@
[…]
@@ 2 Simulation Study
Our experimental design was informed by a combination of related simulation studies, theoretical considerations, and computational constraints. The selected range of sample sizes is consistent with those commonly used in similar studies. However, we include sample sizes of 4 and 8, which are rarely examined. Although such small samples are uncommon in most statistical practice, they may hold practical relevance in fields where data is limited (for example, gene expression studies). For sample sizes of 256 and above, we anticipate that all methods will perform well from a practical standpoint, though these larger sample sizes can also become computationally prohibitive.
The selected range of endpoints includes the 5% and 95% confidence levels and enables the construction of symmetric two-sided confidence intervals at 95% and 90% levels by combining the 0.025 and 0.975, or 0.05 and 0.95 endpoints. These levels are the most commonly used in practice and in almost all simulation studies. However, our choice of endpoints also permits the exploration of other confidence levels. It is important to note, from an evaluation standpoint, that considering one-sided coverage is crucial, as relying solely on two-sided coverage can be misleading (see Section 2.2 for details).
The selected statistical functionals include the most commonly used functionals (mean, median, standard deviation, and correlation) and two extreme percentiles, which are less commonly used in practice but are known to pose challenges in quantifying uncertainty. A notable omission is the inclusion of model coefficients, such as those from the family of generalized linear mixed-effects models. However, for many of these models, and for most more complex models, the bootstrap is not only a viable alternative but the only option for quantifying uncertainty.
We selected the data-generating processes to cover the entire spectrum of difficulty, which is primarily characterized by the skewness and kurtosis of the underlying distribution. A notable omission in our study is the consideration of categorical data, specifically the estimation of proportions in binary data. Addressing categorical data would require a different experimental setup, as our chosen data-generating processes are incompatible with proportion estimation, and the selected statistical functionals do not apply to categorical data. We delegate the analysis of categorical data to future work. However, note that in all but the smallest sample sizes and most extreme proportions, the mean and corresponding quantification of uncertainty will perform well.
We used $n_{\text{rep}}=10000$ replications for every experiment, to limit coverage standard error to $0.005$ in the worst case, and $B=\{10,100,1000\}$ bootstrap replications. More bootstrap replications is better, so we do not consider $B$ a dimension of the experiment. We only report results for $B=1000$, but we would reach the same conclusions with $B=100$. For $B=10$ the performance of bootstrap methods is noticeably worse. Recommendations [3, 8, 38] and choices of $B$ in simulation studies also suggest that $B=1000$ is enough.
@@ 2.1 Methods
As a baseline for comparison, we include the CIs produced by the methods that are most commonly used in practice for that statistical functional. For the mean we use the t-based CIs from the commonly used t-test (t-test). For the median we use CIs from the Wilcoxon signed rank test (wilcoxon[39]. For standard deviation, we use the chi-squared CIs (chi-sq[40]. For Pearson correlation we use Fisher CIs (fisher) [41]. For quantiles we use parametric CIs based on normal assumption (q-par), non-parametric CI (q-nonpar) (see [7] for both), and the Maritz-Jarrett method (m-j) [42].
The bootstrap procedure can be divided into two primary steps. First, bootstrap sampling to generate the bootstrap distribution, which serves as an approximation of the sampling distribution of the functional of interest. And second, applying a bootstrap method to construct a confidence interval.
Bootstrap sampling can be further categorized into parametric and non-parametric approaches. The parametric bootstrap assumes a specific distribution for the underlying population, $F$, and estimates the associated parameters from the observed data, $X=(x_{1},x_{2},\dots,x_{n})$. In contrast, the non-parametric bootstrap infers properties of $F$ by resampling directly from $X$ without imposing distributional assumptions. Given that there are $n^{n}$ possible samples from resampling, we typically restrict the process to $B$ independent bootstrap samples to maintain computational feasibility. This yields $B$ bootstrap samples $X^{*}_{1},X^{*}_{2},\dots,X^{*}_{B}$ and the bootstrap distribution of the parameter $\hat{\theta}^{*}=(\hat{\theta}^{*}_{1},\hat{\theta}^{*}_{2},\dots,\hat{\theta}^{*}_{B})$.
The assumptions underlying parametric bootstrap methods restrict their applicability or require additional user input, making them less suitable as a universal approach. Therefore, our experiments focus exclusively on non-parametric bootstrap methods. Below, we provide a brief overview of the bootstrap methods used for CI construction, along with references for further details. The source code for our implementation is also available (see Section 3).
2.1.1 Percentile Bootstrap (PB)
The percentile method is the original method proposed by Efron in [43] (for details, also see [38, chap. 13]). Even though multiple improvements have been made since, it is still one of the most popular bootstrap methods. The percentile CI for confidence level $\alpha$ is obtained by taking the $\alpha$-quantile of the bootstrap distribution:
$\hat{\theta}_{\text{PB}}[\alpha]=\hat{\theta}^{*}_{\alpha}.$
In all of our implementations of methods that use quantiles, we used the median-unbiased version of quantile calculation, recommended in [44].
2.1.2 Standard Bootstrap (B-n)
The standard method, sometimes also called the normal method, assumes that the bootstrap distribution is normal [38, chap. 13]:
$\hat{\theta}_{\text{B-n}}[\alpha]=\hat{\theta}+\hat{\sigma}z_{\alpha},$
where $\hat{\theta}$ is the plug-in estimator of the functional, $\hat{\sigma}$ is the standard deviation estimate from the bootstrap distribution and $z_{\alpha}$ is the z-score.
2.1.3 Basic Bootstrap (BB)
In the basic bootstrap [38, chap. 13.4], sometimes called the reverse percentile method, we replace the observed bootstrap distribution, $\theta^{*}$ with $W^{*}=\theta^{*}-\hat{\theta}$. This results in
$\hat{\theta}_{\text{BB}}[\alpha]=2\hat{\theta}-\hat{\theta}^{*}_{1-\alpha}.$
Davison and Hinkley [3] show that it provides an accurate confidence interval for the sample median, but it can have a substantial coverage error because of errors in quantile calculation of $W^{*}$. It can also give us invalid parameter values, when there are constraints on $\theta$.
2.1.4 Smoothed Bootstrap (SB)
The smoothed bootstrap [3] gets its name from smoothing the bootstrap distribution. We implement smoothing with a normal kernel centered on 0. We determine kernel size with
$h=0.9\min\big(\hat{\sigma},\frac{\text{IQR}}{1.34}\big),$
where IQR is the inter-quartile range of the bootstrap distribution, respectively. The CI estimate is then obtained by taking the $\alpha$ quantile of the smoothed bootstrap distribution $\tilde{\theta}^{*}$:
$\hat{\theta}_{\text{SB}}[\alpha]=\tilde{\theta}^{*}_{\alpha}.$
2.1.5 Bias Corrected Bootstrap (BC)
The bias corrected bootstrap corrects the bias of the percentile CI [38, chap. 14]. The CI estimate is:
$\hat{\theta}_{\text{BC}}[\alpha]$ $=\hat{\theta}^{*}_{\alpha_{BC}},$
$\alpha_{\text{BC}}$ $=\Phi\big(2\Phi^{-1}(\hat{b})+z_{\alpha}\big),$
where $\Phi$ is the standard normal CDF $\hat{b}$ is the bias, calculated as the percentage of values from the bootstrap distribution that are lower than the value of the functional on the data.
2.1.6 Bias Corrected and Accelerated Bootstrap (BCa)
The bias corrected and accelerated bootstrap [38, chap. 14] further corrects the $BC$ interval by computing acceleration $a$, which accounts for the skewness of the bootstrap distribution:
$\hat{\theta}_{BCa}[\alpha]$ $=\hat{\theta}^{*}_{\alpha_{\text{BCa}}},$
$\alpha_{\text{BCa}}$ $=\Phi\Big(\Phi^{-1}(b)+\frac{\Phi^{-1}(\hat{b})+z_{\alpha}}{1+\hat{a}(\Phi^{-1}(\hat{b})+z_{\alpha})}\Big),$
where $\hat{a}$ is the leave-one-out jackknife approximation of the acceleration constant.
2.1.7 Studentized Bootstrap (B-t)
The studentized bootstrap [38, chap. 14], also known as bootstrap-t, generalizes the Student’s t method, using the distribution of $T=(\hat{\theta}-\theta)/\hat{\sigma}$ to estimate the CI:
$\hat{\theta}_{\text{B-t}}[\alpha]=\hat{\theta}-\hat{\sigma}T_{1-\alpha}.$
But since the distribution of $T$ is not known, we need to approximate its percentiles from the bootstrap distribution. We do that by defining $T^{*}=(\hat{\theta}^{*}-\hat{\theta})/\hat{\sigma}^{*}$, where $\hat{\sigma}^{*}$ is obtained by doing another inner bootstrap sampling on each of the outer samples.
2.1.8 Double Bootstrap (DB)
The double bootstrap [45, chap. 3.11] corrects bias with another level of inner of bootstraps. We repeat the bootstrap procedure on each of the bootstrap samples to calculate the the percentage of times that the inner bootstrap functional is smaller than on the original sample. We want to take such a limit that $P\{\hat{\theta}\in(-\infty,\hat{\theta}_{double}[\alpha])\}=\alpha$, which is why we need to select the $\alpha$-th quantile of biases $\hat{b}^{*}$ for the adjusted level $\alpha_{\text{DB}}$. This leads to
$\hat{\theta}_{\text{DB}}[\alpha]$ $=\hat{\theta}^{*}_{\alpha_{double}},$
$\alpha_{\text{DB}}$ $=\hat{b}^{*}_{\alpha}.$